# Directional Dynamics of the Non-Hermitian Skin Effect

### v1

@@ Abstract
I Introduction
The non-Hermitian skin effect (NHSE)—the dramatic accumulation of all eigenstates at system boundaries under open boundary conditions—has emerged as a defining phenomenon in non-Hermitian physics [4, 28, 10, 2, 15, 37, 11, 16, 18, 27, 44]. Experimental realizations now span photonic lattices [33, 32], topolectrical circuits [12, 14], quantum walks [34], and ultracold atoms [46]. However, existing studies have primarily focused on static properties—spectra, topological invariants, and eigenstate distributions [24, 28, 17, 13, 26, 39]—while the dynamical consequences remain comparatively unexplored [30, 36, 35, 20, 40, 29, 43, 41, 19]. A central open question is how the non-reciprocity parameter $\gamma$ affects information propagation.
@@ I Introduction
Quantum Liang information flow (QLIF) [38, 21, 22] provides precisely such a measure: $T_{B\to A}=dS_{A}/dt-dS_{A\!\not\!B}/dt$, where $S_{A\!\not\!B}$ denotes entropy evolution with subsystem $B$ frozen. Unlike correlations, QLIF is inherently directional ($T_{B\to A}\neq T_{A\to B}$), making it ideally suited for non-reciprocal systems [9]. Here we apply QLIF to the non-Hermitian SSH model [31] with non-reciprocal hopping $t_{1}\pm\gamma$. We discover: (i) a “scissors effect” where $\Delta_{\mathbb{T}}=\mathbb{T}_{R\to L}-\mathbb{T}_{L\to R}$ varies linearly with $\gamma$ for small $\gamma$; (ii) non-monotonic dependence on skin length $\xi$; (iii) velocity ordering $v_{\mathrm{eff}}(\gamma<0)>v_{\mathrm{eff}}(0)>v_{\mathrm{eff}}(\gamma>0)$ revealing NHSE-induced blocking; and (iv) three distinct temporal regimes. These results establish the first quantitative link between skin localization and directional information dynamics.
II Model
@@
III Quantum Liang Information Flow
The quantum Liang information flow (QLIF) [38, 21, 22, 9] quantifies the causal influence of subsystem $B$ on $A$ through a freezing operation. Let $S_{A}(t)=-\mathrm{Tr}[\rho_{A}\ln\rho_{A}]$ be the von Neumann entropy of $A$. We construct a modified Hamiltonian $H_{\!\not\!B}$ with all couplings to $B$ removed, and define the cumulative QLIF as
@@ V Discussion
QLIF overcomes limitations of conventional measures: unlike symmetric correlation functions, it captures the asymmetry $\mathbb{T}_{R\to L}\neq\mathbb{T}_{L\to R}$ characteristic of non-reciprocal dynamics. While Lieb-Robinson bounds [23, 6] rely on quasiparticle propagation [7]—problematic for non-Hermitian systems with biorthogonal eigenstates [16]—QLIF operates directly at the level of entropy production. Our results extend recent Hermitian applications [9] to the non-Hermitian regime.
Several directions remain open: interactions may modify the competition between NHSE localization and many-body delocalization; the interplay with nontrivial topology near $t_{2}=t_{1}$ [19] merits investigation. Experimental verification appears feasible in photonic lattices [33, 32] and topolectrical circuits [14, 12]. Extensions to many-body systems and higher-order skin effects [41] would further illuminate directional dynamics.
VI Conclusion
@@ Acknowledgements.
Appendix A Model Details and Light-Cone Velocity
A.1 Bulk Spectrum and Topological Phases
Under periodic boundary conditions, the Bloch Hamiltonian reads $\mathcal{H}(k)=\bm{d}(k)\cdot\bm{\sigma}$, where
$d_{x}(k)=t_{1}+t_{2}\cos k,\quad d_{y}(k)=t_{2}\sin k+i\gamma.$ (6)
For $\gamma=0$, the dispersion is
$E(k)=\sqrt{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos k},$ (7)
with gap $\Delta=2|t_{1}-t_{2}|$ and winding number $\nu=1$ ($\nu=0$) for $t_{2}<t_{1}$ ($t_{2}>t_{1}$[31].
Table 1 summarizes NHSE parameters for representative $\gamma$ values.
Table 1: NHSE parameters for $t_{1}=1$, $L=42$ sites.
$\gamma$ $r$ $\xi$ Direction IPR
0.4-0.4 1.53 2.4 Right 0.30
0.2-0.2 1.22 4.9 Right 0.12
0 1.00 $\infty$ 0.024
+0.2+0.2 0.82 4.9 Left 0.12
+0.4+0.4 0.65 2.4 Left 0.30
A.2 Derivation of Light-Cone Velocity
Starting from the dispersion relation Eq. (7), the quasiparticle group velocity is
$v_{g}(k)=\frac{\partial E}{\partial k}=-\frac{t_{1}t_{2}\sin k}{E(k)}.$ (8)
To find the maximum of $|v_{g}(k)|$, we analyze
$|v_{g}(k)|=\frac{t_{1}t_{2}|\sin k|}{\sqrt{t_{1}^{2}+t_{2}^{2}+2t_{1}t_{2}\cos k}}.$ (9)
At the critical point $t_{1}=t_{2}\equiv t$, the dispersion simplifies to
$E(k)=t\sqrt{2+2\cos k}=2t\left|\cos\frac{k}{2}\right|,$ (10)
and the group velocity becomes
$|v_{g}(k)|=\frac{t^{2}\cdot 2|\sin(k/2)\cos(k/2)|}{2t|\cos(k/2)|}=t\left|\sin\frac{k}{2}\right|,$ (11)
which is maximized at $k=\pi$, yielding $v_{\mathrm{max}}=t$.
For general parameters, the maximum group velocity is given by
$v_{\mathrm{max}}=\min(t_{1},t_{2}).$ (12)
To prove this, assume $t_{2}\leq t_{1}$ and set $v_{g}(k^{*})=t_{2}$. This gives $t_{1}\sin k^{*}=E(k^{*})$, which after squaring and simplifying yields $(t_{1}\cos k^{*}+t_{2})^{2}=0$, hence $\cos k^{*}=-t_{2}/t_{1}$. This solution exists when $t_{2}\leq t_{1}$, and the corresponding velocity is exactly $v_{g}(k^{*})=t_{2}=\min(t_{1},t_{2})$. This result is analogous to the transverse-field Ising model, where $v_{\mathrm{max}}=\min(J,h)$ with $J$ the exchange coupling and $h$ the transverse field [9].
Table 2 compares the exact numerical values with the approximate formula for $t_{1}=1$.
Table 2: Comparison of exact maximum group velocity with the approximate formula $v_{\mathrm{max}}\approx\min(t_{1},t_{2})$ for $t_{1}=1$.
$t_{2}$ Phase $v_{\mathrm{max}}$ (exact) $\min(t_{1},t_{2})$
0.6 Topological 0.60 0.6
0.8 Topological 0.80 0.8
1.0 Critical 1.00 1.0
1.2 Trivial 1.00 1.0
1.4 Trivial 1.00 1.0
A.3 Lieb-Robinson Bound
The Lieb-Robinson theorem [23, 6] provides a rigorous upper bound on information propagation in lattice systems with local interactions. For a one-dimensional chain with nearest-neighbor hopping of strength $J$, the Lieb-Robinson bound takes the form
$\|[A_{x}(t),B_{0}]\|\leq C\|A\|\|B\|e^{-a(|x|-v_{\text{LR}}|t|)},$ (13)
where $v_{\text{LR}}$ is the Lieb-Robinson velocity and $a$ is a decay constant. For tight-binding models, a standard estimate gives $v_{\text{LR}}\sim 2J_{\max}$ where $J_{\max}$ is the largest hopping amplitude [6]. For the Hermitian SSH model ($\gamma=0$) with hopping amplitudes $t_{1}$ and $t_{2}$, we use
$v_{\text{LR}}=2\max(t_{1},t_{2}).$ (14)
For our parameters ($t_{1}=1$, $t_{2}=0.5$), this yields $v_{\text{LR}}=2.0$.
The bound guarantees exponential suppression outside the light cone but does not forbid small leakage. This explains why onset times in Fig. 6 can appear slightly “superluminal”—we detect the exponentially small signal beyond the strict light cone.
Note that $v_{\text{LR}}\geq v_{\max}$ always holds: the Lieb-Robinson bound is a rigorous upper limit, while the maximum group velocity $v_{\max}=\min(t_{1},t_{2})$ derived from the dispersion relation is typically tighter. For $t_{1}=1$, $t_{2}=0.5$: $v_{\max}=0.5$ (per unit cell) versus $v_{\text{LR}}=2.0$.
A.4 NHSE-Induced Velocity Ordering
The onset time $t^{*}$ (earliest time with $|\mathbb{T}|>10^{-6}$) reveals $\gamma$-dependent velocity ordering:
This reflects NHSE blocking: $\gamma>0$ localizes eigenstates toward the left boundary, impeding rightward propagation. At large distances ($d\gtrsim 10$), this blocking becomes dramatic as information must propagate against the skin direction.
Appendix B Numerical Methods
B.1 Exact Diagonalization
We numerically diagonalize the non-Hermitian SSH Hamiltonian using standard linear algebra routines. For non-Hermitian matrices, we compute both right and left eigenvectors, though for QLIF calculations only the right eigenvectors are needed.
B.2 Time Evolution and QLIF Implementation
For a single-particle state $|\psi\rangle=\sum_{j}\psi_{j}c_{j}^{\dagger}|0\rangle$, the site entropy reduces to binary entropy:
$S_{A}=-|\psi_{A}|^{2}\ln|\psi_{A}|^{2}-(1-|\psi_{A}|^{2})\ln(1-|\psi_{A}|^{2}).$ (16)
The freezing operation sets all matrix elements connected to site $B$ to zero:
$(H_{\!\not\!B})_{ij}=\begin{cases}0&\text{if }i=B\text{ or }j=B,\\ H_{ij}&\text{otherwise}.\end{cases}$ (17)
Time evolution is performed via exact diagonalization:
$|\psi(t)\rangle=e^{-iHt}|\psi_{0}\rangle=\sum_{n}e^{-iE_{n}t}|n\rangle\langle n|\psi_{0}\rangle,$ (18)
where $\{E_{n},|n\rangle\}$ are eigenvalues and right eigenvectors. For non-Hermitian $H$, biorthogonal normalization is employed.
B.3 Inverse Participation Ratio (IPR)
The inverse participation ratio (IPR) is a standard measure of localization [8] and has been widely used to characterize the skin effect in non-Hermitian systems [42, 25]. It is defined as
$\text{IPR}=\sum_{j}|\psi_{j}|^{4}\bigg/\left(\sum_{j}|\psi_{j}|^{2}\right)^{2},$ (19)
ranging from $1/L$ (extended) to $1$ (fully localized).
Here we derive the IPR for eigenstates exhibiting the non-Hermitian skin effect.
For a skin-effect eigenstate localized at the left boundary, the probability amplitude decays exponentially:
$|\psi_{n}|^{2}=A\,r^{2(n-1)},\quad n=1,2,\ldots,N,$ (20)
where $n$ labels the unit cell, $r=|\beta|$ is the skin parameter [Eq. (LABEL:eq:r)], and $A$ is the normalization constant.
Normalization.
From $\sum_{n=1}^{N}|\psi_{n}|^{2}=1$:
$A\sum_{n=1}^{N}r^{2(n-1)}=A\frac{1-r^{2N}}{1-r^{2}}=1,$ (21)
giving
$A=\frac{1-r^{2}}{1-r^{2N}}.$ (22)
IPR calculation.
The IPR is
IPR $=\sum_{n=1}^{N}|\psi_{n}|^{4}=A^{2}\sum_{n=1}^{N}r^{4(n-1)}$
$=A^{2}\frac{1-r^{4N}}{1-r^{4}}=\left(\frac{1-r^{2}}{1-r^{2N}}\right)^{2}\frac{1-r^{4N}}{(1-r^{2})(1+r^{2})}$
$=\frac{(1-r^{2})(1-r^{4N})}{(1-r^{2N})^{2}(1+r^{2})}.$ (23)
Using the identity $1-r^{4N}=(1-r^{2N})(1+r^{2N})$:
$\text{IPR}_{\text{th}}=\frac{1-r^{2}}{1+r^{2}}\cdot\frac{1+r^{2N}}{1-r^{2N}}.$ (24)
Limiting cases.
  • Strong localization ($r\ll 1$, NN large): $\text{IPR}\to(1-r^{2})/(1+r^{2})\to 1$.
  • Extended states ($r=1$): Using L’Hôpital’s rule, $\text{IPR}\to 1/N$, the expected value for uniform distribution.
  • Finite-size correction: For $r<1$ and large $N$, $\text{IPR}\approx(1-r^{2})/(1+r^{2})$.
The IPR values listed in Table 1 are obtained from direct numerical diagonalization, which includes finite-size corrections and deviations from the idealized exponential decay near boundaries. For the Hermitian case ($r=1$), the table correctly shows $\text{IPR}=1/L\approx 0.024$ for $L=42$ sites.
B.4 Oscillation Analysis
Late-time QLIF oscillations have period $T_{\text{osc}}\approx 2.2$ independent of system size $L$, ruling out finite-size boundary reflections. The period is determined by intrinsic energy-level spacing: $T_{\text{osc}}\sim 2\pi/\Delta E$ where $\Delta E\approx 2.9$ matches the bulk bandwidth.
The NHSE protects oscillation amplitude: comparing $L=20$ to $L=120$, the Hermitian amplitude decays by $\sim 3\times$ while the non-Hermitian ($\gamma=0.3$) amplitude decays by only $\sim 1.3\times$, because skin-localized eigenstates maintain high local probability density.
B.5 Sublattice Configuration for QLIF Measurements
The SSH model possesses a bipartite structure with alternating $\alpha$ (even sites) and $\beta$ (odd sites) sublattices. When measuring the directional asymmetry of QLIF, it is essential that all relevant sites—the initial excitation site $j_{0}$ and the left/right observation sites—reside on the same sublattice.
If the observation sites are placed on a different sublattice from the initial state, the different hopping structure (intracell vs. intercell) introduces an intrinsic asymmetry even in the Hermitian limit ($\gamma=0$). Specifically, for $\gamma=0$ with mixed-sublattice configuration:
$\mathbb{T}_{R\to L}(t)\neq\mathbb{T}_{L\to R}(t),$ (25)
with differences reaching $|\Delta_{\mathbb{T}}|\sim 0.07$ at $t=10$. This spurious asymmetry is entirely a consequence of the sublattice mismatch, not the non-Hermitian physics.
By contrast, when all sites are on the same sublattice (e.g., all on $\alpha$), the Hermitian case exhibits perfect symmetry:
$\mathbb{T}_{R\to L}(t)=\mathbb{T}_{L\to R}(t)\quad(\gamma=0,\text{ same sublattice}),$ (26)
with numerical deviations at the level of machine precision ($\sim 10^{-15}$).
For the results presented in this work, we use the “same-sublattice” configuration: the initial state is placed on site $j_{0}=20$ ($\alpha$ sublattice), with observation sites at $j_{0}\pm d$ where $d$ is an even integer to ensure all sites remain on the $\alpha$ sublattice. The standard configuration uses $d=6$, corresponding to a distance of 3 unit cells.

### v2

@@ Abstract
pacs:
03.65.Vf, 11.30.Er, 05.60.Gg, 73.20.Fz
1 Introduction
The non-Hermitian skin effect (NHSE)—the dramatic accumulation of all eigenstates at system boundaries under open boundary conditions—has emerged as a defining phenomenon in non-Hermitian physics [4, 28, 10, 2, 15, 42, 11, 16, 18, 27, 48]. Experimental realizations now span photonic lattices [37, 34], topolectrical circuits [12, 14], quantum walks [20, 38], and ultracold atoms [50]. While the relationship between non-reciprocal couplings and skin accumulation direction is well established in one-dimensional models [42, 27, 48], recent works have shown that this correspondence can be nontrivially modified or even reversed in more general settings [41, 29, 30]. However, existing studies have primarily focused on static properties—spectra, topological invariants, and eigenstate distributions [35, 25, 28, 17, 13, 26, 44]—while the dynamical consequences remain comparatively unexplored [36, 32, 40, 39, 21, 45, 31, 47, 46, 19]. A central open question is how the non-reciprocity parameter $\gamma$ affects information propagation.
@@ 1 Introduction
Quantum Liang information flow (QLIF) [43, 22, 23] provides precisely such a measure: $T_{B\to A}=dS_{A}/dt-dS_{A\!\not\!B}/dt$, where $S_{A\!\not\!B}$ denotes entropy evolution with subsystem $B$ frozen [8]. Unlike correlations, QLIF is inherently directional ($T_{B\to A}\neq T_{A\to B}$), making it ideally suited for non-reciprocal systems [9]. Here we apply QLIF to the non-Hermitian SSH model [33] with non-reciprocal hopping $t_{1}\pm\gamma$. We discover: (i) a “scissors effect” where $\Delta_{\mathbb{T}}=\mathbb{T}_{R\to L}-\mathbb{T}_{L\to R}$ varies linearly with $\gamma$ for small $\gamma$; (ii) non-monotonic dependence on skin length $\xi$; (iii) velocity ordering $v_{\mathrm{eff}}(\gamma<0)>v_{\mathrm{eff}}(0)>v_{\mathrm{eff}}(\gamma>0)$ revealing NHSE-induced blocking; and (iv) three distinct temporal regimes. These results establish the first quantitative link between skin localization and directional information dynamics.
2 Model
@@
3 Quantum Liang Information Flow
[…]
@@ 5 Discussion
QLIF overcomes limitations of conventional measures: unlike symmetric correlation functions, it captures the asymmetry $\mathbb{T}_{R\to L}\neq\mathbb{T}_{L\to R}$ characteristic of non-reciprocal dynamics. While Lieb-Robinson bounds [24, 6] rely on quasiparticle propagation [7]—problematic for non-Hermitian systems with biorthogonal eigenstates [16]—QLIF operates directly at the level of entropy production [8]. Our results extend recent Hermitian applications [9] to the non-Hermitian regime.
We note that our analysis relies on the standard correspondence between the non-reciprocity parameter $\gamma$ and the skin accumulation direction in the one-dimensional SSH model [42]. However, in more general settings—such as higher-dimensional lattices, photon-mediated interactions, or waveguide-coupled arrays—the direction of eigenstate accumulation does not always follow that of the asymmetric couplings, and skin effect reversal can occur [41, 29, 30]. Investigating how QLIF signatures—particularly the sign rule Eq. (4) and the velocity ordering Eq. (5)—are modified in systems exhibiting such reversed NHSE constitutes an intriguing direction for future work.
Several other directions remain open: interactions may modify the competition between NHSE localization and many-body delocalization; the interplay with nontrivial topology near $t_{2}=t_{1}$ [19] merits investigation. Experimental verification appears feasible in photonic lattices [37, 34] and topolectrical circuits [14, 12]. Extensions to many-body systems and higher-order skin effects [46] would further illuminate directional dynamics.
6 Conclusion